OpenStudy (anonymous):

simplify sqrt(90) A.(9)sqrt(10) B.(10)sqrt(9) c.(3)sqrt(10) d.(10)sqrt(3)

OpenStudy (anonymous):

let's find factors of 90 that are perfect squares :)

OpenStudy (anonymous):

so its b

OpenStudy (anonymous):

\[\sqrt{90}=\sqrt{9*10}\] 9 would be a perfect square so we should be able to simplify and bring it out from underneath the radical

OpenStudy (anonymous):

not going to be b

OpenStudy (anonymous):

a

OpenStudy (anonymous):

not quite, remember that when we simplify it we are looking for what number times itself equals 9 \[\sqrt{90}=\sqrt{9*10}=\sqrt{3*3*10}=3\sqrt{10}\]

OpenStudy (anonymous):

ok can u help mw dis one sqrt(121)a^3b^2 A.11absqrt(a) B.11ab c.121ab 11absqrt11ab

OpenStudy (anonymous):

the best thing to do is break everything down into factors, then any factor appearing twice (like the 3 before) gets to come out of the radical \[\sqrt{121a ^{3}b ^{2}}\]

OpenStudy (anonymous):

\[=\sqrt{11*11*a*a*a*b*b}\] can you simplify from there?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well 11 is twice so it comes out, same with a and b, but there is one a left over and it will have to stay under the radical. \[11ab \sqrt{a}\]

OpenStudy (anonymous):

Use the following to check your work:\[\left\{3 \sqrt{10},9 \sqrt{10},30,3 \sqrt{10},10 \sqrt{3}\right\} \]